OpenStudy (anonymous):

Find the equation in polar form: -7x^(2)-3xy-7y^(2)= -2

OpenStudy (anonymous):

-7 x^2-3 x y-7 y^2+2 = 0

OpenStudy (anonymous):

Simply put rcosθ in place of x & rsinθ in place of y & solve it we have 7*r^2*(cosθ)^2+3*r*Sinθ*r*Cosθ+7*r^2*(Sinθ)^2=2 Now realising that (cosθ)^2+(Sinθ)^2=1 we have 7*r^2((cosθ)^2+(Sinθ)^2)+3*r^2*Sinθ*Cosθ=2 & r^2*(7+3*Sinθ*Cosθ)=2 & r=sqrt(2/(7+3*Sinθ*Cosθ)) This is the answer. More details about this curve is here http://www.wolframalpha.com/input/?i=+-7x%5E%282%29-3xy-7y%5E%282%29%3D+-2

OpenStudy (anonymous):